Water flows from a tap at rate of 10 meters/seconds into a cylindrical tank whose diameter is 2 meter and whose height is 4.9 meters. If the inner radius of the tap is 3.5 cm, find the time required to fill this tank in minutes and seconds. (Take  Π =22/7)
 
                                  

Grade:10

1 Answers

Khimraj
6 years ago
Volume of tank = πr2h = (22/7)(2/2)2(4.9) =15.4m3
Let it takes time t seconds then
(volume of water through tap per second)*t = 15.4m3
(22/7)(3.5/100)2(10)*t = 15.4m3
(0.0385m3/s)*t = 15.4m3
t = 400 seconds = 6 minutes and 40 seconds.
Hope it clears. If ypu like anser then please approve it.